# Long and Short Guidance in Score identity Distillation for One-Step Text-to-Image Generation

### v3

@@ 2.3 Long and Short Guidance
Previous works equipped with a fake score network $f_{\psi}$ typically only consider adding CFG when evaluating the pretrained score network $f_{\phi}$, such as in DMD (Yin et al., 2023). In the context of SiD, this corresponds to setting $\kappa_{1}=\kappa_{2}=\kappa_{3}=1$ and $\kappa_{4}>1$. In this paper, we discover that a broad spectrum of combinations of $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$, and $\kappa_{4}$ can all significantly enhance performance compared to not using any CFG at all. These settings are found to lead to different balances of minimizing FID, which reflects how well the generated data match the training data in distribution, and maximizing the CLIP score (Radford et al., 2021), which reflects how well the generated images follow the textual guidance.
This flexibility to accommodate various CFG combinations expands the design space for SD distillation, balancing generation quality and text adherence. However, given the vastness of the search space, we are motivated to develop strategies that constrain the scope of exploration.
@@
Short strategy: Weakening CFG of the fake score network $f_{\psi}$. With the availability of a fake score network $f_{\psi}$, we have developed a new strategy to enhance SiD’s text-to-image generation capabilities by reducing the CFG of $f_{\psi}$ in SiD. An exemplary configuration is setting $\kappa_{1}=\kappa_{4}=1$ while allowing $\kappa_{2}=\kappa_{3}$ to vary within $(0,1)$. The underlying idea is that by diminishing $f_{\psi}$’s capacity to detect generations aligned with ${\boldsymbol{c}}$ through a reduced CFG scale ($10<\kappa_{2}<$ 1), the student generator is incentivized to produce images that better align with the textual guidance to compensate for this reduction. We will present experiments to assess the efficacy and limitations of this strategy.
Long and short CFGs. We introduce an innovative strategy termed long-short guidance (LSG), where we enhance CFG during the training of $f_{\psi}$ by setting $\kappa_{1}>1$, and maintain CFGs on $f_{\psi}$ and $f_{\phi}$ during the training of $G_{\theta}$ at the same or a lower scale by setting $1\leq\kappa_{2}=\kappa_{3}=\kappa_{4}\leq\kappa_{1}$. The logic behind this approach is that enhancing the CFG of $f_{\psi}$ during training effectively reduces the CFG at evaluation time when used at the same or a reduced level. To our knowledge, we are the first to incorporate CFG into the training of the fake score network. LSG has been found to achieve a better balance between FID and CLIP scores, optimizing performance across these metrics.
@@ 3 Experiments
Figure 3: Left (Long CFG of the true score network): This plot illustrates the progression of FID and CLIP scores, each influenced by different CFGs applied to the true score network. The $\kappa$’s not specified in the legends are set as 1. FID scores are plotted on the primary y-axis, while CLIP scores are displayed on the secondary y-axis in corresponding line styles but with slight transparency. Together, these lines demonstrate how various CFGs impact model performance. Right (Short CFG of the fake score network): Analogous plot to the left where the CFGs of the fake score network are either weakened during evaluation or enhanced during training.
@@
We summarize the parameter settings, such as batch size, learning rate, and optimizer configurations, in Table 3 in Appendix C. Unless specified in ablation studies, the settings are uniformly applied across all guidance strategies. We present the details of our method in Algorithm 1, where we generate an image in one step and generalize the setting in DMD and SiD to set $\omega(t)$ as
@@
where $C$ is the total number of pixels in $\boldsymbol{x}_{g}$. We initialize both ${\psi}$ and ${\theta}$ using the pretrained ${\phi}$ from SD. We have performed an ablation study that chooses $t_{init}\in\{354,550,625,675,800,900,999\}$ and found that the model’s performance under FP32 optimization is not sensitive to these choices. To align with the setting in SiD, we set $t_{init}=625$, which would result in $\sigma_{t_{init}}/a_{t_{init}}=2.5$ under the DDPM schedule used to train SD1.5. We conduct a comprehensive study to evaluate the performance of SiD using the proposed LSG for distilling SD1.5. Additionally, we apply the same LSG scales to distill SD2.1-base, further assessing the adaptability and effectiveness of our approach across different model versions.
We consider a standard setting that utilizes the Aesthetics6+ prompt (Cherti et al., 2023) for training and evaluates performance by computing zero-shot FID on the COCO-2014 validation set. We adhere to the standard protocol by generating 30k images to compare with the 40,504 images in the COCO-2014 validation set for calculating zero-shot FID. Additionally, we employ the ViT-g-14-laion2b_s12b_b42k encoder (Ilharco et al., 2021; Cherti et al., 2023) to compute the CLIP score (Radford et al., 2021). The FID and CLIP scores presented in the figures are calculated using randomly sampled prompts from the COCO-2014 validation set. When reporting the FID and CLIP results of SiD in Table 1, we follow DMD (Yin et al., 2023) to use the exact evaluation code [^1] https://github.com/mingukkang/GigaGAN/tree/main/evaluation provided by GigaGAN (Kang et al., 2023), where a pre-defined list of 30k text prompts selected from the COCO-2014 validation set is used to generate the 30k images, which are used for computing FID and CLIP scores with images from the validation set as the reference.
@@
Table 2: Comparison of HPSv2 score and Precision/Recall on the COCO-2014 validation set. The HPSv2 scores of ADD (SD-Turbo) are produced based on the publicly available model checkpoint. The HPSv2 scores of the other baselines are quoted from SwiftBrush (Nguyen and Tran, 2024). The Precision and Recall on COCO-2014 are obtained using the 30K images generated by the corresponding model checkpoints. The SwiftBrush checkpoint was provided upon request by the authors.
Teacher Student Human Preference Score v2 $\uparrow$ COCO-2014 Precision and Recall $\uparrow$
@@
SD2.1-base SwiftBrush (Nguyen and Tran, 2024) 26.91 27.21 26.32 26.37 0.47 0.46
@@ 3.1 Long and/or short guidance strategies
Short the CFG of the fake score network. For the short strategy, we set $\kappa_{1}=\kappa_{4}=1$ and explore $\kappa_{2}=\kappa_{3}\in\{0.5,0.125\}$. As illustrated in the right panel of Figure 4, this approach delivers competitive performance, with $\kappa_{2}$ dictating the balance between FID and CLIP scores. However, compared to the long strategy, this configuration generally produces inferior results, as suggested by lower CLIP scores when FIDs are controlled to similar levels. From Figure 4, it’s also evident that when No CFG is employed, where $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=1$, the results are underwhelming, underscoring the importance of injecting CFGs into SiD to distill text-to-image diffusion models.
LSG: Long and Short classifier-free Guidance. Below, we explore how to effectively integrate the long and short strategies to enhance text guidance in diffusion distillation. Initially, we discovered an alternative approach to shorting the CFG of the fake score network $f_{\psi}$: by amplifying the CFG during the training of $f_{\psi}$. This approach represents the “simplest” form of LSG strategy. Specifically, we set $\kappa_{1}$ to 2 or 3, while keeping $\kappa_{2}=\kappa_{3}=\kappa_{4}$ at 1. As illustrated in the right panel of Figure 4, setting $\kappa_{1}=3$ (green lines) proves as effective as the short strategy with $\kappa_{2}=\kappa_{3}=0.125$ (red lines), and $\kappa_{1}=2$ (black lines) outperforms the short strategy with $\kappa_{2}=\kappa_{3}=0.5$ (blue lines) and is comparable to the long strategy with $\kappa_{4}=3$ (red lines in the left panel). These findings validate this “simplest” LSG as an effective guidance strategy in text-guided diffusion distillation.
@@ 3.2 Ablation study
We investigate the impact of extended training durations under two different guidance scales, variations in batch size, and the selection of training prompts. Initially, we doubled the number of fake images used to train the generator from 10M to 20M and monitored the evolution of the FID and CLIP scores. From the left panel of Figure 5, we observe continuous improvements in both FID and CLIP scores for an LSG of 1.5 when training is extended beyond 10M fake images, and sustained enhancements in CLIP scores for an LSG of 2.0. Notably, by doubling the training input from 10M to 20M fake images, the FID under LSG 1.5 decreased from 8.71 to a record low of 8.15 among diffusion distillation methods.
Additionally, we assessed the effects of batch size by considering two additional settings: a batch size of 8192 with a learning rate of 4e-6, a configuration used in SiD to distill the EDM model pretrained on ImageNet 64x64 (Zhou et al., 2024), and a batch size of 64 with a learning rate of 1e-6. The middle panel of Figure 5 demonstrates that while there are initial differences in the convergence speed in terms of the number of fake images processed, the performances eventually converge to similar levels. We note that while smaller batch sizes may seem to converge faster, they require more time to process the same number of images. This is due to more frequent model parameter updates, higher communication costs between GPUs, and additional overheads typically associated with smaller batch sizes.
@@ 3.3 Quantitative and qualitative evaluations
SD1.5 distilled with SiD-LSG ($\kappa=1.5$): FID $=$ 8.15, CLIP $=$ 0.304 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
SD2.1-base distilled with SiD-LSG ($\kappa=4.5$): FID = 16.54, CLIP = 0.322 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
Figure 7: Visual comparison of two SiD-LSG models: one preferred for FID and the other for CLIP. All images are generated from the same text prompt: “A distinguished older gentleman in a vintage study, surrounded by books and dim lighting, his face marked by wisdom and time. 8K, hyper-realistic, cinematic, post-production." The model with a lower guidance scale of $\kappa=1.5$, which achieves a record-low one-step-generation FID of 8.15 and a competitive CLIP score of 0.304, produces images that are more diverse but align less closely with specific text details, such as “dim lighting.” Conversely, the model with a higher guidance scale of $\kappa=4.5$, achieving a high CLIP score of 0.322 and noted for state-of-the-art human preference scores (HPSv2) as shown in Table 2, presents a relatively high FID of 16.54, indicating less diversity but superior text alignment and visual quality.
SD1.5 distilled with SiD-LSG ($\kappa=1.5$): FID $=$ 8.15, CLIP $=$ 0.304 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
SD2.1-base distilled with SiD-LSG ($\kappa=4.5$): FID = 16.54, CLIP = 0.322 Refer to caption   Refer to caption   Refer to caption   Refer to caption   Refer to caption
Figure 8: Analogous to Figure 8, this plot compares two rows of images generated using two distinct guidance scales, both conditioning on the same text prompt: “poster art for the collection of the asian woman, in the style of gloomy, dark orange and white, dynamic anime, realistic watercolors, nintencore, weathercore, mysterious realism –ar 69:128 –s 750 –niji 5".
We present comprehensive results from prior studies across various experimental settings, including both one-step and multi-step generation methods. When evaluation results are available in existing literature (Yin et al., 2023; Liu et al., 2023; Kang et al., 2023; Nguyen and Tran, 2024), we directly cite them; otherwise, if model checkpoints are accessible, either publicly or provided upon request by the authors, we utilize the evaluation code from GigaGAN to produce the reported results. For our SiD-LSG, we select $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}\in\{1.5,2.0,3.0,4.5\}$.
For comparisons of FID and CLIP scores, the results are detailed in Table 1. Among all one-step distillation methods, our approach notably excels in zero-shot text-conditioned image generation on the COCO-2014 dataset, as reflected by both FID-30K and CLIP scores. Specifically, with the guidance scale set as 2, our method attains FID scores as low as 9.56 and 10.97, and a CLIP score above 0.31 and around 0.32, using SD 1.5 and 2.1-base as the pretrained backbones, respectively. Notably, by setting the guidance scale to 1.5 and doubling the training time, our data-free method achieves a record-low FID of 8.15, along with a CLIP score of 0.304, when distilling SD1.5. These results remain highly competitive when compared to other generative approaches, such as autoregressive models and GANs, and are even comparable to previous multi-step diffusion-based sampling methods. Analyzing different $\kappa$ values, we observe a trade-off between FID and CLIP scores: smaller $\kappa$ values generally yield better FID metrics, while larger values enhance CLIP scores, aligning with past findings on the impact of guidance scale.
Beyond FID and CLIP scores, we also assess Precision and Recall (Kynkäänniemi et al., 2019) as well as Human Preference Score (HPSv2) (Wu et al., 2023), which are presented in Table 2. We reuse the same 30k images from previous evaluations for the Precision and Recall calculations. For HPSv2, we follow their established protocol, generating images from 800 text prompts per category. Our SiD-LSG models outperform other baselines in HPSv2 scores across all categories, as well as in Precision and Recall metrics. Notably, with $\kappa=4.5$, SiD-LSG reaches peak performance in distilling SD2.1-base. Regarding Precision and Recall, higher $\kappa$ values lead to improved Precision, while lower values result in better Recall.
@@
To contextualize the numerical differences in metrics such as FID, CLIP, and HPSv2, we present Figures 8 and 8 to illustrate what these differences mean for visual perception: the model with the best FID score excels in diversity, while the model with the best CLIP score stands out in text alignment and aesthetic quality.
@@
For broader impact, please refer to the discussion in Appendix A. For limitations and computational requirements, please refer to a detailed discussion in Appendix D.
4 Conclusion

### v4

@@ 2.3 Long and Short Guidance
Previous works equipped with a fake score network $f_{\psi}$ typically only consider adding CFG when evaluating the pretrained score network $f_{\phi}$, such as in DMD (Yin et al., 2023). In the context of SiD, this corresponds to setting $\kappa_{1}=\kappa_{2}=\kappa_{3}=1$ and $\kappa_{4}>1$. In this paper, we discover that a broad spectrum of combinations of $\kappa_{1}$, $\kappa_{2}$, $\kappa_{3}$, and $\kappa_{4}$ can all significantly enhance performance compared to not using any CFG at all, which means setting $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=1$. These different combinations are found to lead to different balances of minimizing FID, which reflects how well the generated data match the training data in distribution, and maximizing the CLIP score (Radford et al., 2021), which reflects how well the generated images follow the textual guidance. This flexibility to accommodate various CFG combinations expands the design space for SD distillation, balancing generation quality and text adherence. However, given the vastness of the search space, we are motivated to develop strategies that constrain the scope of exploration. We present three such strategies, acknowledging that there may be more effective approaches not explored in this paper.
@@
Short strategy: Weakening CFG of the fake score network $f_{\psi}$. With the availability of a fake score network $f_{\psi}$, we have developed a new strategy to enhance SiD’s T2I generation capabilities by reducing the CFG of $f_{\psi}$ in SiD. An exemplary configuration is setting $\kappa_{1}=\kappa_{4}=1$ while allowing $\kappa_{2}=\kappa_{3}$ to vary within $(0,1)$. The underlying idea is that by diminishing $f_{\psi}$’s capacity to detect generations aligned with ${\boldsymbol{c}}$ through a reduced CFG scale ($10<\kappa_{2}<$ 1), the student generator is incentivized to produce images that better align with the textual guidance to compensate for this reduction. We will present experiments to assess the efficacy and limitations of this strategy.
Long and short CFGs. We introduce an innovative strategy termed long-short guidance (LSG), where we enhance CFG during the training of $f_{\psi}$ by setting $\kappa_{1}>1$, and maintain CFGs on $f_{\psi}$ and $f_{\phi}$ during the training of $G_{\theta}$ at the same or a lower scale by setting $1\leq\kappa_{2}=\kappa_{3}=\kappa_{4}\leq\kappa_{1}$. The logic behind LSG is that enhancing the CFG of $f_{\psi}$ during training effectively reduces its CFG at evaluation time when used at the same or a reduced level. To our knowledge, we are the first to incorporate CFG into the training of the fake score network $f_{\psi}$. LSG has been shown to effectively balance reducing FID scores while increasing CLIP scores, and will therefore be used as the default method.
We refer the reader to Appendix C for further discussion on the long and short CFG strategies.
@@ 3 Experiments
Figure 3: Left (Long CFG of the true score network): This plot illustrates the gradual decline in FID and the corresponding rise in CLIP scores, each influenced by different CFGs applied to the true score network. $\kappa$ values not specified in the legend are set to 1. FID scores are plotted on the primary y-axis, while CLIP scores are displayed on the secondary y-axis in corresponding line styles but with slight transparency. Together, these lines demonstrate how various CFGs impact model performance. Right (No CFG; Short CFG of the fake score network with $\kappa_{2}=\kappa_{3}\in(0,1)$; a simple form of LSG that sets $\kappa_{1}>1$): Analogous plot to the left where the CFGs of the fake score network are not applied, weakened during evaluation, or enhanced during training.
@@
We summarize the parameter settings, such as batch size, learning rate, and optimizer configurations, in Table 4 in Appendix D. Unless specified in ablation studies, the settings are uniformly applied across all guidance strategies. We present the details of our method in Algorithm 1, where we generate an image in one step and generalize the setting in DMD and SiD to set $\omega(t)$ as
@@
where $C$ is the total number of pixels in $\boldsymbol{x}_{g}$. We initialize both ${\psi}$ and ${\theta}$ using the pretrained ${\phi}$ from SD. We have tested $t_{init}\in\{354,550,625,675,800,900,999\}$ and found that the model’s performance under FP32 optimization is not sensitive to these choices. To align with the setting in SiD, we set $t_{init}=625$, which would result in $\sigma_{t_{init}}/a_{t_{init}}=2.5$ under the DDPM schedule used to train SD1.5. We conduct a comprehensive study to evaluate the performance of SiD using the proposed LSG for distilling SD1.5. Additionally, we apply the same LSG scales to distill SD2.1-base, further assessing the adaptability and effectiveness of our approach across different model versions.
We consider a standard setting that utilizes the Aesthetics6+ prompt (Cherti et al., 2023) for training and evaluates performance by computing zero-shot FID on the COCO-2014 validation set. We adhere to the standard protocol by generating 30k images to compare with the 40,504 images in the COCO-2014 validation set for calculating zero-shot FID. Additionally, we employ the ViT-g-14-laion2b_s12b_b42k encoder (Ilharco et al., 2021; Cherti et al., 2023) to compute the CLIP score (Radford et al., 2021). The FID and CLIP scores presented in the figures are calculated using randomly sampled prompts from the COCO-2014 validation set. When reporting the FID and CLIP results of SiD in Table 1, we use the exact evaluation code [^1] https://github.com/mingukkang/GigaGAN/tree/main/evaluation provided by GigaGAN (Kang et al., 2023), where a pre-defined list of 30k text prompts selected from the COCO-2014 validation set is used to generate the 30k images, which are used for computing FID and CLIP scores with images from the validation set as the reference.
@@
Table 2: Comparison of HPSv2 score and Precision/Recall on the COCO-2014 validation set. The HPSv2 scores of ADD (SD-Turbo) are produced based on the publicly available model checkpoint. The HPSv2 scores of the other baselines are quoted from SwiftBrush (Nguyen & Tran, 2024). The Precision and Recall on COCO-2014 are obtained using the 30K images generated by the corresponding model checkpoints.
Teacher Student Human Preference Score v2 $\uparrow$ COCO-2014 Precision & Recall $\uparrow$
@@
SD2.1-base SwiftBrush (Nguyen & Tran, 2024) 26.91 27.21 26.32 26.37 0.47 0.46
@@ 3.1 Long and/or short guidance strategies
No CFG. Figure 4 demonstrates that without CFG, where $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}=1$, results are underwhelming, highlighting the need for CFG in SiD to distill T2I diffusion models.
@@
Short the CFG of the fake score network. For the short strategy, we set $\kappa_{1}=\kappa_{4}=1$ and explore $\kappa_{2}=\kappa_{3}\in\{0.5,0.125\}$. As illustrated in the right panel of Figure 4, this approach delivers competitive performance, with $\kappa_{2}$ dictating the balance between FID and CLIP scores. However, compared to the long strategy, this configuration generally produces inferior results, as suggested by lower CLIP scores when FIDs are controlled to similar levels.
LSG: Long and Short classifier-free Guidance. Below, we explore how to effectively integrate the long and short strategies to enhance text guidance in diffusion distillation. Initially, we discovered a “simplest” form of LSG strategy by amplifying the CFG during the training of $f_{\psi}$. Specifically, we set $\kappa_{1}$ to 2 or 3, while keeping $\kappa_{2}=\kappa_{3}=\kappa_{4}$ at 1. As illustrated in the right panel of Figure 4, setting $\kappa_{1}=3$ (green lines) proves as effective as the short strategy with $\kappa_{2}=\kappa_{3}=0.125$ (red lines), and $\kappa_{1}=2$ (black lines) outperforms the short strategy with $\kappa_{2}=\kappa_{3}=0.5$ (blue lines) and is comparable to the long strategy with $\kappa_{4}=3$ (red lines in the left panel). These findings validate this “simplest” LSG as an effective guidance strategy in text-guided diffusion distillation.
@@ 3.2 Ablation study
We investigate the impact of extended training durations under two different guidance scales, variations in batch size, and the selection of training prompts. Initially, we doubled the number of fake images used to train the generator from 10M to 20M and monitored the evolution of the FID and CLIP scores. From the left panel of Figure 5, we observe continuous improvements in both FID and CLIP scores for an LSG of 1.5 when training is extended beyond 10M fake images, and sustained enhancements in CLIP scores for an LSG of 2.0. Notably, by doubling the training input from 10M to 20M fake images, the FID under LSG 1.5 decreased from 8.71 to a record low of 8.15 among diffusion distillation methods in the data-free setting.
@@
We assessed the effects of batch size by considering two additional settings: a batch size of 8192 with a learning rate of 4e-6, a configuration used in SiD to distill the EDM model pretrained on ImageNet 64x64 (Zhou et al., 2024), and a batch size of 64 with a learning rate of 1e-6. The middle panel of Figure 5 demonstrates that while there are initial differences in the convergence speed in terms of the number of fake images processed, the performances eventually converge to similar levels. We note that while smaller batch sizes may seem to converge faster, they require more time to process the same number of images. This is due to more frequent model parameter updates, higher communication costs between GPUs, and additional overheads typically associated with smaller batch sizes.
@@ 3.3 Quantitative and qualitative evaluations
We present comprehensive results from prior studies across various experimental settings, including both one-step and multi-step generation methods. When evaluation results are available in existing literature (Yin et al., 2023; Liu et al., 2023; Kang et al., 2023; Nguyen & Tran, 2024), we directly cite them; otherwise, if model checkpoints are accessible, either publicly or provided upon request by the authors, we utilize the evaluation code from GigaGAN to produce the reported results. For our SiD-LSG, we select $\kappa_{1}=\kappa_{2}=\kappa_{3}=\kappa_{4}\in\{1.5,2.0,3.0,4.5\}$.
For comparisons of FID and CLIP scores, the results are detailed in Table 1. Among all one-step distillation methods, our approach notably excels in zero-shot text-conditioned image generation on the COCO-2014 dataset, as reflected by both FID-30K and CLIP scores. Specifically, with the guidance scale set as 2, our method attains FID scores as low as 9.56 and 10.97, and a CLIP score above 0.31 and around 0.32, using SD 1.5 and 2.1-base as the pretrained backbones, respectively. Notably, by setting the guidance scale to 1.5 and doubling the training time, our method achieves a record-low data-free FID of 8.15, along with a CLIP score of 0.304, when distilling SD1.5. These results remain highly competitive when compared to other generative approaches, such as autoregressive models and GANs, and are even comparable to previous multi-step diffusion-based sampling methods. Analyzing different $\kappa$ values, we observe a trade-off between FID and CLIP scores: smaller $\kappa$ values generally yield better FID metrics, while larger values enhance CLIP scores, aligning with past findings on the impact of guidance scale.
Beyond FID and CLIP scores, we also assess Precision and Recall (Kynkäänniemi et al., 2019) as well as Human Preference Score (HPSv2) (Wu et al., 2023), which are presented in Table 2. We reuse the same 30k images from previous evaluations for the Precision and Recall calculations. For HPSv2, we follow their established protocol, generating images from 800 text prompts per category. Except for ADD trained with real data and adversarial loss, our SiD-LSG models outperform other baselines in HPSv2 scores across all categories, as well as in Precision and Recall metrics. Notably, with $\kappa=4.5$, SiD-LSG reaches peak performance in distilling SD2.1-base. Regarding Precision and Recall, higher $\kappa$ values lead to improved Precision, while lower values result in better Recall.
@@
To contextualize the numerical differences in metrics such as FID, CLIP, and HPSv2, we present Figures 8 and 8 to illustrate their implications for visual perception: the model with the best FID excels in diversity, while the model with the best CLIP stands out in text alignment and aesthetic quality.
@@
For broader impact, please refer to the discussion in Appendix A. For limitations and computational requirements, please refer to a detailed discussion in Appendix E.
4 Conclusion and Future Work